OpenStudy (anonymous):

What is the solution set of the following system of equations? 5x - 3y - 10 = 0 x + y - 7 = 0 A{(31/8, 25/8)} B{(25/8, 31/8)} C{(45/2, 31/2)} D{(31/2, 45/2)}

OpenStudy (anonymous):

i think A is write: 5x-3y-10=0 => y=(5/3)x-10/3 x+y-7=0 => y=-x+7 {y=(5/3)x-10/3 {y=-x+7 multiply second equation in (-1) => (-1)(y=-x+7) so y=(5/3)x-10/3 -y=x-7 add two equation (add left side with left side & rught side with right side) then: 0=(8/3)x-31/3 (8/3)x=31/3 x=31/8 y=-x+7 => (x=31/8)=>y=31/8 =>y=25/8 so : x=31/8 & y=25/8